\begin{thebibliography}{}

\bibitem{kaysh14}
Cantarini N., Kac V.,
Classification of linearly compact simple Jordan and generalized Poisson superalgebras,
J. Algebra, 313 (2007), 100--124.

\bibitem{kacan}
Cantarini N., Kac V.,
Classification of linearly compact simple rigid superalgebras,
Int. Mat. Res. Not., 17 (2010), 3341--3393.

\bibitem{Chapa01}
Chapoton F.,
Un endofoncteur de la categorie des operades, Dialgebras and related operads, 105--110,
Lect. Notes in Math., 1763, Springer, 2001.


    \bibitem{dedkov89}
    Dedkov A.,
    Some properties of quasi-associative and quasi-alternative algebras,
    Siberian Math. J., 30 (1989), 3, 479--483.


\bibitem{Dorofeev}
Dorofeev G. V.,
On nilpotency of right alternative rings,
Algebra and Logic,  9 (1970), 3, 302--305.

\bibitem{Dzhuma01}
Dzhumadil'daev A., Abdykassymova S.,
Leibniz algebras in characteristic p,
C. R. Acad. Sci. Paris Ser. I Math., 332 (2001), 12, 1047--1052.

\bibitem{Dzhuma07}
Dzhumadil'daev A.,
Zinbiel algebras over a q-commutator, J. Math. Sci. (N. Y.), 144 (2007),  2, 3909--3925.

\bibitem{Dzhuma11}
Dzhumadil'daev A., Ismailov N., A., Tulenbaev K.,
Free bicommutative algebras, Serdica Math. J., 37 (2011), 1, 25--44.


\bibitem{Dzhuma14}
Dzhumadil'daev A., Ismailov N.,
$S_n$- and $GL_n$-module structures on free Novikov algebras,
Journal of Algebra,  416 (2014),  287--313.


\bibitem{Fialc12}
 Fialowski A., Khudoyberdiyev A., Omirov B.,
A characterization of nilpotent Leibniz algebras,
Algebras and Repr. Theory,  16 (2013), 5, 1489--1505.

    \bibitem{Kantor72}
    Kantor I.,
    Certain generalizations of Jordan algebras,  Trudy Sem. Vektor. Tenzor. Anal.,
    16 (1972), 407--499. [in Russian]

    \bibitem{Kantor90}
    Kantor I., The universal conservative algebra, Siberian Math. J., 31 (1990),  3, 388--395.




\bibitem{KLP}
Kaygorodov I., Lopatin A., Popov Yu.,
Conservative algebras of 2-dimensional algebras,
Linear Algebra and its Applications,  486 (2015),  255--274.

\bibitem{KV15}
 Kaygorodov I., Volkov Yu., 
 Conservative algebras of $2$-dimensional algebras, II,
Comm. Alg. (2017).   

\bibitem{KP}
Kaygorodov I., Popov Yu.,
Alternative algebras admitting derivations with invertible values and invertible derivations,
Izvestiya: Mathematics, 78 (2014), 5, 922--936.




    \bibitem{shestakov_kuzm}
    Kuzmin E., Shestakov I.,
    {\it Nonassociative structures,}
    Itogi Nauki i Tekhniki: Sovremennye Problemy Mat.:
    Fundamental'nye Napravleniya, vol. 57, VINITI, Moscow 1990, pp. 179--266;
    English transl., A.I. Kostrikin and I. R. Shafarevich (eds.),
    Encyclopaedia Math. Sci., vol. 57, Algebra VI, Springer-Verlag, London--New York 1995, pp. 199--280.

\bibitem{Loday07}
Loday J.-L., On the algebra of quasi-shuffles,
Manuscripta mathematica, 123(1) (2007), 79--93.

\bibitem{Loday08}
Loday J.-L.,
Generalized bialgebras and triples of operads,
Asterisque, 320 (2008), x+116 p.

\bibitem{shestakov93}
Makar-Limanov L., Shestakov I.,
Polynomial and Poisson dependence in free Poisson algebras and free Poisson fields, 
J. Algebra, 349 (2012),
372--379.


\bibitem{kolesn08}
Pozhidaev A. P., Dialgebras and related triple systems,
Sib. Math. J., 49 (2008),  4, 696--708


\bibitem{xu97}
Zakharov A.,
Novikov-Poisson algebras and superalgebras of Jordan brackets,
Comm. Algebra, 42 (2014),  5, 2285--2298.


\bibitem{rings}
Zhevlakov K., Slinko A., Shestakov I., Shirshov A.,
Rings that are nearly associative, Pure and Applied Mathematics, 104. Academic Press, Inc.
 [Harcourt Brace Jovanovich, Publishers],
New York-London, 1982. xi+371 pp.

\bibitem{zinbiel}
Zinbiel G. W.,
Encyclopedia of types of algebras 2010,
Operads and universal algebra, 217–297,
Nankai Ser. Pure Appl. Math. Theoret. Phys., 9, World Sci. Publ., Hackensack, NJ, 2012.

    \end{thebibliography}

\documentclass[10pt,oneside,reqno]{amsart}
\usepackage[utf8]{inputenc}
\usepackage[english]{babel}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{mathrsfs}

\newtheorem{Th}{Theorem}
\newtheorem{Lem}[Th]{Lemma}

\begin{document}
    \sloppy

    {\huge On The Kantor Product }

    \

   {\bf Ivan Kaygorodov\\
}
    \medskip

    Universidade Federal do ABC, CMCC, Santo Andr\'{e}, Brazil.

 \


\

\

{\bf Abstract. }
We study the algebra of bilinear multiplications of an $n$-dimensional vector space.
In particular, we study the Kantor product of some well-known
(associative, Lie, alternative, Novikov and some other) multiplications.



    \section{Introduction}

Kantor introduced the class of conservative algebras in \cite{Kantor72}.
This class includes some well-known algebras: associative, Jordan, Lie, Leibniz and Zinbiel
\cite{KLP}.
In the theory of conservative algebras of great importance is the conservative algebra $U(n)$ \cite{Kantor90}.
In the theory of Lie algebras  $U(n)$  plays a role analogous to the role of $\mathfrak{gl}_n.$
The space of the algebra $U(n)$ is the space of all bilinear multiplications on
the $n$-dimensional space $V_n$.
To define  the operation of multiplication $[ \ , \ ]$ in the algebra $U(n)$
we fix a vector $u \in V_n$ and
for two multiplications $A, B \in U(n)$ and two elements $x,y \in V_n$ we  set
\begin{eqnarray}\label{Kanpro}
x*y=[A,B](x,y) = A(u,B(x,y))-B(A(u,x),y)-B(x,A(u,y)).
\end{eqnarray}
Some properties of the algebra $U(2)$ were studied in \cite{KLP,KV15}.
We say that the product of two multiplications on $n$-dimensional vector space
defined by (\ref{Kanpro}), is the left Kantor product of these multiplications.
In a similar way we can define the right Kantor product and obtain similar results. 
We assume that the Kantor product is the left Kantor product.
The Kantor product of a multiplication  $\cdot$ by itself is {\it the Kantor square} of $\cdot$ and it is denoted by
 $[\cdot, \cdot].$
It gives us a map $K$ from any variety $V$ of algebras to some class $K(V).$ 
 
The Kantor square of a multiplication $\cdot$ can be rewritten (see \cite{Kantor72})
 as the product of the left multiplication $L_u$
and the multiplication $\cdot,$ as $[L_u, \cdot],$ where
$$[L_u, \cdot](x,y) = u\cdot(x\cdot y)- (u\cdot x)\cdot y - x \cdot (u\cdot y)=[\cdot, \cdot](x,y).$$
The multiplication $[L_u, \cdot]$ plays an important role in the definition of a (left) conservative algebras
 \cite{kacan,Kantor72}.
We recall that an algebra $A$ with a multiplication $\cdot$ is called a (left) conservative algebra
if and only if there exist a new multiplication $*$  such that
$$[L_a,[L_b, \cdot]]=- [L_{a*b},\cdot].$$

The main aim of this paper is to study the properties of the Kantor product of multiplications.
One of the central questions studied in this paper
is the following:

{\it\bf Question.} What identities does the class of algebras $K(V)$ satisfy if we know the identities of $V$?

We give some particularly  answer of this question for 
associative, (anti)commutative, Perm, Lie, Leibniz, Zinbiel, left-commutative, bicommutative,
Novikov, alternative, quasi-associative and quasi-alternative algebras;
 we also describe the Kantor product of multiplications
in associative dialgebras, duplicial, dual duplicial, $As^{(2)}$,
Poisson, generalized Poisson and Novikov-Poisson algebras.
Finally,  we study the Kantor square in some special cases;
in particular,
the associative algebras with identities, nilpotent and right-nilpotent algebras,
associative algebras isomorphic to its Kantor square;
and discuss the coincidence of derivations and automorphisms of the algebra and its Kantor square.
Here we can to formulate

{\it\bf Open problem.} Is $K(V)$ a variety of algebras for some variety $V$?


    \section{The Kantor square}

\

In this section we leave technical and trivial proofs of lemmas.
We are using the standard notation:
$$(a,b,c)_*=(a*b)*c-a*(b*c), (a,b,c)=(ab)c-a(bc); $$
$$[a,b]_*=a*b-b*a, [a,b]=ab-ba;$$
$$\circlearrowright_{a,b,c} [f(a,b,c)]=f(a,b,c)+f(b,c,a)+f(c,a,b).$$

\subsection{Associative algebras.}
The variety of {\it associative algebras} is defined by the identity
$$(xy)z=x(yz).$$

    \begin{Lem}\label{thass} Let $(A;\cdot)$ be an associative algebra.
    Then $(A;[\cdot, \cdot])$ is an associative algebra.
    \end{Lem}

\

\subsection{(Anti)commutative algebras.}
The variety of {\it  $($anti$)$commutative algebras} is defined by the identity
$$xy=\epsilon yx,$$ where $\epsilon=1$ in the commutative case
 and $\epsilon =-1$ in the  anticommutative case.

    \begin{Lem}\label{anticom} Let $(A;\cdot)$ be an $($anti$)$commutative algebra.
    Then $(A;[\cdot, \cdot])$ is an $($anti$)$commutative algebra.
    \end{Lem}

\

\subsection{Perm algebras.}
The variety of {\it Perm  algebras} (see, for example, \cite{Chapa01}) is defined by the identity
$$(xy)z=x(yz)=x(zy).$$

    \begin{Lem} Let $(A;\cdot)$ be a Perm algebra.
    Then $(A;[\cdot, \cdot])$ is a Perm  algebra.
    \end{Lem}

\

\subsection{Lie algebras.}
The variety of {\it Lie algebras} is defined by the identities
$$xy=-yx, (xy)z+(yz)x+(zx)y=0.$$

    \begin{Lem}Let $(A;\cdot)$ be a Lie algebra.
    Then $[\cdot, \cdot]=0.$
    \end{Lem}

\

\subsection{Leibniz algebras.}
The variety of (left) {\it Leibniz algebras} (see, for example, \cite{Dzhuma01})
 is defined by the identity
$$x(yz)=(xy)z+y(xz).$$

    \begin{Lem}\label{leibniz}Let $(A;\cdot)$ be a $($left$)$ Leibniz algebra.
    Then $[\cdot, \cdot]=0.$
    \end{Lem}

\

\subsection{Left-commutative algebras.}
The variety of {\it  left-commutative algebras} (see, for example, \cite{shestakov_kuzm})
 includes commutative-associative, bicommutative, Novikov, Zinbiel algebras and some other.
This variety is defined by the identity
$$x(yz)=y(xz).$$

    \begin{Lem}\label{thleftcom} Let $(A;\cdot)$ be a left-commutative algebra.
    Then $(A;[\cdot, \cdot])$ is a left-commutative algebra.
    \end{Lem}

\

\subsection{Bicommutative algebras.}
The variety of {\it bicommutative algebras} (see, for example, \cite{Dzhuma11})
 is defined by the identities
$$x(yz)=y(xz), (xy)z=(xz)y.$$

    \begin{Lem} Let $(A;\cdot)$ be a bicommutative algebra.
    Then $(A;[\cdot, \cdot])$ is an associative-commutative algebra.
    \end{Lem}

\

\subsection{Zinbiel algebras.}
The variety of (left) {\it  Zinbiel algebras} (see, for example, \cite{Dzhuma07})
is defined by the identity
$$x(yz)=(xy)z+(yx)z.$$

    \begin{Lem} Let $(A;\cdot)$ be a $($left$)$ Zinbiel algebra.
    Then $(A;[\cdot, \cdot])$ is a $($left$)$ Zinbiel algebra.
    \end{Lem}

\

\subsection{Novikov algebras.}
The variety of (left) {\it Novikov algebras} (see, for example, \cite{Dzhuma14})
 is defined by the identities
$$x(yz)=y(xz), (x,y,z)=(x,z,y).$$

    \begin{Lem}Let $(A;\cdot)$  be a $($left$)$ Novikov algebra.
    Then  $(A;[\cdot, \cdot])$ is a $($left$)$ Novikov algebra.
    \end{Lem}

\

\subsection{Alternative algebras.}
The variety of {\it alternative algebras} (see, for example, \cite{KP})
is defined by the identities
\begin{eqnarray}\label{altern}
x^2y=x(xy), xy^2=(xy)y.
\end{eqnarray}
It is also well known (see, for example, \cite{rings})
that an  alternative algebra is flexible:  $$(xy)x=x(yx);$$ 
and satisfies the Moufang identities:
$$x(yzy)=((xy)z)y, (yzy)x=y(z(yx)), (xy)(zx)=x(yz);$$
and the following identities hold:
$$(x,y,z)=-(y,x,z), (x,y,z)=-(x,z,y),$$
$$(x,xy,z)=(x,y,z)x,(x,yx,z)=x(x,y,z).$$

The main example of a non-associative alternative algebra is
 a Cayley --- Dickson algebra ${\bf C}$ \cite{rings}.
Let $F$ be a field of characteristic $\neq 2.$ 
It is an algebra ${\bf C}$ with the basis $e_0 = 1, e_1,\dots,e_7$ 
and the following multiplication table:
\begin{equation}
\label{table}
\begin{array}{|c||c|c|c|c|c|c|c|} \hline
1 &  e_1 & e_2 & e_3 & e_4 & e_5 & e_6 & e_7 \\
\hline \hline e_1 & \alpha \cdot 1 &  e_3  &  \alpha e_2 & e_5 &  \alpha e_4  & -e_7 &  -  \alpha e_6 \\
\hline
 e_2  & -e_3 & \beta \cdot 1 & - \beta e_1 & e_6 & e_7 &  \beta  e_4 &  \beta e_5 \\
\hline
 e_3  & - \alpha e_2 &  \beta e_1 & -\alpha \beta \cdot 1 & e_7 &  \alpha e_6 &  -\beta e_5 &
 -\alpha \beta  e_4 \\
\hline
 e_4 & -e_5 & -e_6 & -e_7 & \gamma \cdot 1 & - \gamma e_1 & -\gamma e_2 &  -\gamma e_3 \\
\hline
 e_5 & - \alpha e_4 & -e_7 & -\alpha  e_6 & \gamma e_1 & -\alpha\gamma \cdot 1  &  \gamma  e_3 &
  -\alpha \gamma e_2 \\
\hline
 e_6 & e_7 & -\beta e_4 &  \beta e_5 & \gamma e_2 & -\gamma e_3 &  -\beta \gamma \cdot 1 &
 -\beta\gamma\cdot e_1 \\
\hline
 e_7 & \alpha  e_6 & -\beta e_5 &  \alpha \beta e_4 &  \gamma e_3 &  -\alpha \gamma e_2 &
 \beta \gamma e_1 & \alpha\beta\gamma \cdot 1 \\
\hline
\end{array}
\end{equation}

    \begin{Th}\label{teoremaalt}Let $(A;\cdot)$  be an alternative algebra.
    Then  $(A;[\cdot, \cdot])$ is a flexible algebra. Furthermore,\\
    1$)$ $(A;[\cdot, \cdot])$ is an alternative algebra if and only if $A$ satisfies the identity
    \begin{eqnarray}\label{altass}(x,u,(x,u,y))=0;\end{eqnarray}
    2$)$ $(A;[\cdot, \cdot])$ is a noncommutative Jordan algebra if and only if $A$
    satisfies the identity
    $$[L_uL_xL_uL_x, R_uR_x] = [L_{xuxu}, R_{ux}];$$
    3$)$ $(A;[\cdot, \cdot])$ is a Jordan algebra if $(A, \cdot)$ is a commutative alternative algebra$;$\\
    4$)$ $({\bf C};[\cdot, \cdot])$ is alternative for a Cayley --- Dickson algebra ${\bf C}$,
    if and only if $u= u_0\cdot 1.$
    \end{Th}

{\bf Proof.}
It is easy to see that
$$a * b = u(ab)-(ua)b-a(ub)= (au)b- 2a(ub).$$

Now, we can see
$$ (x*y)*x-x*(y*x)= ((xu)y-2x(uy))*x- x*((yu)x-2y(ux))=$$
$$(((xu)y)u)x-2((x(uy))u)x-2((xu)y)(ux)+4(x(uy))(ux)-$$
$$(xu)((yu)x)+2(xu)(y(ux))+2x(u((yu)x))-4x(u(y(ux)))=$$
$$x(uyu)x-2((x,u,y)u)x-2x(uyu)x-2x(uyu)x+2((xu)y,u,x)+4x(uyu)x-$$
$$x(uyu)x-2(xy)(y,u,x)+2x(uyu)x+2x(uyu)x+2x(u(y,u,x))-4x(uyu)x=$$
$$2[((xu)y,u,x)-(x,u,x(uy))-((xy)(y,u,x)-x(u(y,u,x)))]=$$
$$2[((x,u,y),u,x)-((x,u,y),u,x)]=0.$$

It follows that $(A,*)$ is a flexible algebra.

$1)$ It is easy to see that a flexible algebra is alternative if and only if it satisfies the first identity from (\ref{altern}).
We have
$$(x*x)*y - x*(x*y)=$$
$$-(xuxu)y+2(xux)(uy)
-(xu)((xu)y)+2x(u((xu)y))+2(xu)(x(uy))-4x(u(x(uy)))=$$
$$-2(xu)^2y-2x(u(x(uy)))+2x(u((xu)y))+2(xu)(x(uy))=(x,u,(x,u,y)).$$
Now, the multiplication $[\cdot, \cdot]$ is alternative if and only if $(x,u,(x,u,y))=0.$

$2)$ Note that an algebra $B$ is a {\it  non-commutative Jordan algebra}
if and only if it is flexible and it satisfies the Jordan identity:
$(x^2,y,x)=0.$

Obviously,
$$((x*x)*y)*x-(x*x)*(y*x)=$$
$$(-(xuxu)y+2(xux)(uy))*x + (xux)*((yu)x-2y(ux))=$$
$$-(((xuxu)y)u)x+2((xuxu)y)(ux)+2(((xux)(uy))u)x-4((xux)(uy))(ux)+$$
$$(xuxu)((yu)x)-2(xux)(u((yu)x))-2(xuxu)(y(ux))+4(xux)(u(y(ux)))=$$
$$2(((xuxu)y)(ux)-(xuxu)(y(ux)) +((x(u(x(uy))))u)x-x(u(x(u((yu)x)))).$$

Now, $[\cdot, \cdot]$ is a noncommutative Jordan multiplication  if and only if
$$[L_xL_uL_xL_u, R_uR_x] = [L_{xuxu}, R_{ux}].$$

$3)$ It is easy to see that
if $A$ is a commutative alternative algebra then we have
$$((xuxu)y)(ux)-(xuxu)(y(ux)) +(((xux)(uy))u)x-(xux)(u((yu)x))=$$
$$(xu)((xu)((xu)y)))- (xu)((xu)((xu)y)))+
x(u(x(u(x(uy)))))-x(u(x(u(x(uy)))))=0.$$
It follows that $[\cdot, \cdot]$ is  non-commutative Jordan 
and from Theorem \ref{anticom} we infer that $[\cdot, \cdot]$ is  Jordan.

$4)$ If $({\bf C},*)$  is an alternative algebra for every $u$
then $A$ satisfies (\ref{altass}).
Note that for the elements
$e_{i_1},e_{i_2},e_{i_3},$
where $e_{i_k}e_{i_l}\neq \epsilon e_{i_m}$ (where $\epsilon$ is some element from the ground field),
 we have
$$(e_{i_1}, e_{i_2}, (e_{i_1},e_{i_2},e_{i_3}))= 2 (e_{i_1})^2(e_{i_2})^2 e_{i_3}.$$
Such triple $(i_1,i_2,i_3)$ we  call a g-triple.
It is easy to see that if $(i,j,k)$ is not a g-triple then
 the subalgebra generated by $e_i, e_j, e_k$ is a two-generated subalgebra,
 and by the Artin theorem
 this subalgebra is  associative, i. e., $(e_i,e_j, (e_i,e_j,e_k))=0.$
 Now,
 for the element $u=u_0\cdot 1 + u_1 e_1 +\ldots+u_7 e_7$
 we have $(e_i, u, (e_i, u, e_j))=0$ if and only if
$\sum\limits_{\mbox{k, (i,j,k) is a g-triple}} (u_k e_k)^2 =0.$
It is equivalent to the following system
\begin{equation*}
\begin{array}{ccccccccc}
\alpha u^2_1 &              & -\alpha\beta u^2_3 & + \gamma u^2_4 &                      &
-\beta\gamma u^2_6 &                          & =& 0, \\
\alpha u^2_1 & +\beta u^2_2 &                    &                & -\alpha \beta u_5^2  &
-\beta\gamma u^2_6 &                          & =& 0, \\
\alpha u^2_1 & +\beta u^2_2 &                    & + \gamma u^2_4 &                      &
& +\alpha\beta\gamma u_7^2  & =& 0, \\
\alpha u^2_1 &              & -\alpha\beta u^2_3 &                & -\alpha \beta u_5^2  &
                  & +\alpha\beta\gamma u_7^2  & =& 0, \\
             & +\beta u^2_2 & -\alpha\beta u^2_3 & + \gamma u^2_4 & -\alpha \beta u_5^2  &
                               &                          & =& 0, \\
             & +\beta u^2_2 & -\alpha\beta u^2_3 &                &                      &
             -\beta\gamma u^2_6 & +\alpha\beta\gamma u_7^2  & =& 0, \\
             &              &                    & + \gamma u^2_4 & -\alpha \beta u_5^2  &
             -\beta\gamma u^2_6 & +\alpha\beta\gamma u_7^2  & =& 0.  \\
\end{array}
\end{equation*}

Calculating,  we obtain
$u_1=\sqrt{\beta \gamma} u_7, u_6=\sqrt{-\alpha} u_7, u_2=u_3=u_4=u_5=0.$

Now, from the relation (\ref{altass})
by simple calculations (for example, for $x=e_1+e_2, y=e_1$ and $x=e_2+e_6,y=e_6$)
we can find that $u_7=0$ and $u= u_0\cdot 1.$

The theorem is proved.

\subsection{Quasi-associative algebras.}
{\it Quasi-associative  algebras} (see, for example, \cite{dedkov89})
is defined by the identities
$$(x,y,z)+(y,z,x)+(z,x,y)=0,$$
$$(x,y,z) = \alpha [y,[x,z]],$$
where $\alpha$ is a fixed element in the ground field $F.$
It is known \cite{dedkov89} that an algebra $(A, \cdot)$ is quasi-associative
if and only if there exist an associative algebra
$A$ with the new multiplication, such that for some $\lambda\in F$:
$$x\cdot y = \lambda xy + (1-\lambda) yx.$$

    \begin{Lem} Let $(A;\cdot)$ be a quasi-associative algebra.
    Then $(A;[\cdot, \cdot])$ is a quasi-associative algebra.
    \end{Lem}

\

\subsection{Quasi-alternative algebras.}
{\it Quasi-alternative  algebras} (see, for example, \cite{dedkov89})
is defined by the identities
$$(x,y,x)=0,$$
$$(x,x,y) = \alpha [x,[x,y]],$$
where $\alpha$ is a fixed element from the ground field $F.$
It is known \cite{dedkov89} that an algebra $(A, \cdot)$ is a quasi-alternative algebra 
if and only there exist an alternative algebra $A$ with new multiplication, such that for some $\lambda\in F$:
$$x\cdot y = \lambda xy + (1-\lambda) yx.$$

    \begin{Lem} Let $(A;\cdot)$ be a quasi-alternative algebra.
    Then $(A;[\cdot, \cdot])$ is a flexible algebra.
    \end{Lem}

\

\subsection{Associative dialgebras.}
The variety of {\it associative dialgebras} (see, for example, \cite{kolesn08})
is defined by the identities
$$(x \vdash y) \vdash z = ( x \dashv y)  \vdash z, x \dashv ( y \vdash z) =  x \dashv (y   \dashv z),$$
$$(x \vdash y) \vdash z = x \vdash (y \vdash z), (x \dashv y) \dashv z = x \dashv (y \dashv z),
(x \vdash y) \dashv z = x \vdash (y \dashv z).$$

    \begin{Lem}Let $(A;\vdash,\dashv)$ be an associative dialgebra.
    Then  $(A;[\vdash, \dashv])$  is an associative algebra.
    \end{Lem}

\

\subsection{Duplicial algebras.}

The variety of {\it duplicial algebras} (see, for example, \cite{Loday08}) is defined by the identities
$$(x \prec y) \prec z = x \prec (y \prec z ),$$
$$(x \succ y) \prec z = x \succ (y \prec z),$$
$$(x \succ y) \succ z = x \succ (y \succ z).$$

    \begin{Lem}Let $(A;\prec,\succ)$ be a duplicial algebra.
    Then  $(A;[\succ ,\prec ])$  is an associative algebra.
    \end{Lem}

\subsection{Dual duplicial algebras.}

The variety of {\it dual duplicial algebras} (see, for example, \cite{zinbiel})
 is defined by the identities
$$(x \prec y) \prec z = x \prec (y \prec z ),(x \succ y) \prec z = x \succ (y \prec z),(x \succ y)
\succ z = x \succ (y \succ z),$$
$$ x \prec  (y \succ z) = ( x \prec  y) \succ z =0.$$

    \begin{Lem}Let $(A;\prec,\succ)$ be a dual duplicial algebra.
    Then  $[\succ ,\prec ]=0$ and $(A;[\prec , \succ ])$  is a $2$-nilpotent algebra.
    \end{Lem}

\subsection{$As^{(2)}$-algebras.}

The variety of {\it  $As^{(2)}$-algebras} (see, for example, \cite{zinbiel}) is defined by the identities
$$(x \circ y) \cdot z = x \circ (y \cdot z), (x \cdot y) \circ  z = x \cdot (y \circ z),  $$
$$(x \circ y) \circ z = x \circ (y \circ z), (x \cdot y) \cdot  z = x \cdot (y \cdot z).  $$

    \begin{Lem}Let $(A;\cdot,\circ)$ be a $As^{(2)}$-algebra.
    Then  $(A;[\cdot,\circ ])$ and $(A;[\circ  , \cdot ])$  are associative algebras.
    \end{Lem}

\subsection{Commutative tridendriform algebra.}
The variety of {\it commutative tridendriform algebras} (see, for example, \cite{Loday07})
is defined by the identities
$$x \cdot y  = y \cdot x, (x \cdot y) \cdot z = x \cdot (y \cdot z),$$
$$(x \prec y) \prec z = x \prec (y \prec z ) + x \prec (z \prec y ),$$
$$(x\cdot y)  \prec z=x \cdot (y  \prec z).$$

    \begin{Lem}Let $(A;\cdot,\prec)$ be a commutative tridendriform algebra.
    Then   $(A;[\prec,\cdot ])$ is a commutative algebra and  $(A;[\cdot  , \prec ])$
    is a right Zinbiel algebra.
    \end{Lem}

\subsection{Poisson algebras.}
The variety of {\it Poisson algebras} (see, for example, \cite{shestakov93}) is defined by the identities
$$xy=yx, (xy)z=x(yz), \{xy,z\}=\{x,z\}y+x\{y,z\},$$
$$\{x,y\}=-\{y,x\}, \{\{x,y\},z\}+\{\{y,z\},x\}+\{\{z,x\},y\}=0.$$

    \begin{Th}\label{ThPoisson} Let $(A;\cdot,\{,\})$ be a Poisson algebra.
    Then  $[\{,\}, \cdot]=0$ and  $(A;[\cdot, \{,\}])$ is a Lie  algebra.
    \end{Th}

{\bf Proof.}
In the first case, we have
$$a*b= \{u,ab\} - \{u,a\}b-a\{u,b\}=0.$$

In the second case,
$$a*b= u\{a,b\}-\{ua,b\}-\{a,ub\}= - a\{u,b\}- b\{a,u\}-u\{a,b\}=-b*a.$$
and
$$(a*b)*c+(b*c)*a+(c*a)*b=$$
$$-(a\{u,b\}+\{a,u\}b+\{a,b\}u)*c$$
$$  -(b\{u,c\}+\{b,u\}c+\{b,c\}u)*a$$
$$  -(c\{u,a\}+\{c,u\} a+\{c,a\}u)*b=$$

$$  a\{u,b\}\{u,c\}+\{a\{u,b\},u\}c+\{a\{u,b\},c\}u+$$
$$    \{a,u\}b\{u,c\}+\{\{a,u\}b,u\}c+\{\{a,u\}b,c\}u+$$
$$ \{a,b\}u\{u,c\}+\{\{a,b\}u,u\}c+\{\{a,b\}u,c\}u+$$
$$    b\{u,c\}\{u,a\}+\{b\{u,c\},u\}a+\{b\{u,c\},a\}u+$$
$$ \{b,u\}c\{u,a\}+\{\{b,u\}c,u\}a+\{\{b,u\}c,a\}u+$$
$$    \{b,c\}u\{u,a\}+\{\{b,c\}u,u\}a+\{\{b,c\}u,a\}u+$$
$$ c\{u,a\}\{u,b\}+\{c\{u,a\},u\}b+\{c\{u,a\},b\}u+$$
$$   \{c,u\}a\{u,b\}+\{\{c,u\}a,u\}b+\{\{c,u\}a,b\}u+$$
$$\{c,a\}u\{u,b\}+\{\{c,a\}u,u\}b+\{\{c,a\}u,b\}u=$$

$$a\{ u,b\}\{ u,c\}+\{ u,b\}\{ a,u\}c+ca\{\{ u,b\} ,u\}+au\{ \{ u,b\},c\}+\{ u,b\}\{ a,c\}u+$$
$$\{ a,u\}b\{ u,c\}+\{ \{ a,u\},u\}bc+\{a ,u\}\{ b,u\}c+bu\{\{a ,u\} ,c\}+\{ a,u\}\{ b,c\}u+$$
$$\{ a,b\}u\{ u,c\}+\{ \{ a,b\},u\}uc+\{ a,b\}\{ u,c\}u+uu\{ \{ a,b\},c\}+$$
$$b\{ u,c\}\{ u,a\}+\{ b,u\}\{ u,c\}a+ba\{ \{ u,c\},u\}+\{ \{ u,c\},a\}bu+\{ u,c\}\{ b,a\}u+$$
$$\{ b,u\}c\{ u,a\}+\{ b,u\}\{ c,u\}a+ca\{ \{ b,u\},u\}+\{ b,u\}\{ c,a\}u+\{ \{ b,u\},a\}cu+$$
$$\{ b,c\}u\{ a,u\}+\{ \{ b,c\},u\}ua+\{ b,c\}\{ u,a\}u+uu\{ \{ b,c\},a\}+$$
$$c\{ u,a\}\{ u,b\}+\{ c,u\}\{ u,a\}b+cb\{ \{ u,a\},u\}+\{ c,b\}\{ u,a\}u+cu\{ \{ u,a\},b\}+$$
$$\{ c,u\}a\{ u,b\}+\{ \{ c,u\},u\}ab+\{ c,u\}\{ a,u\}b+\{ c,u\}\{ a,b\}u+\{ \{ c,u\},b\}au+$$
$$\{ c,a\}u\{ u,b\}+\{ \{ c,a\},u\}ub+\{ c,a\}\{ u,b\}u+uu\{ \{ c,a\},b\}=$$

$$  (a\{ u,b\}\{ u,c\}+ \{ c,u\}a\{ u,b\})+
    (\{ u,b\}\{ a,u\}c+ \{a ,u\}\{ b,u\}c)+$$
$$  (ca\{\{ u,b\} ,u\}+ ca\{ \{ b,u\},u\})+
    (\{ u,b\}\{ a,c\}u+ \{ c,a\}\{ u,b\}u)+$$
$$  (\{ a,u\}b\{ u,c\}+ b\{ u,c\}\{ u,a\})+
    (\{ \{ a,u\},u\}bc+ cb\{ \{ u,a\},u\})+$$
$$  (\{ a,u\}\{ b,c\}u+ \{ b,c\}u\{ a,u\})+
    (\{ a,b\}u\{ u,c\}+ \{ c,u\}\{ a,b\}u)+$$
$$  (\{ a,b\}\{ u,c\}u+ \{ b,u\}\{ u,c\}a)+
    (ba\{ \{ u,c\},u\}+ \{ \{ c,u\},u\}ab)+$$
$$  (\{ b,u\}c\{ u,a\}+ \{ c,u\}\{ u,a\}b)+
    (\{ b,u\}\{ c,a\}u+ \{ c,a\}u\{ u,b\})+$$
$$  (\{ b,c\}\{ u,a\}u+ \{ c,b\}\{ u,a\}u)+
    (\{ c,u\}\{ u,a\}b+ \{ c,u\}\{ a,u\}b)+$$
$$  [au\{ \{ u,b\},c\}+au\{ \{ b,c\},u\}+au\{ \{ c,u\},b\}]+$$
$$  [bu\{\{a ,u\} ,c\}+bu\{ \{ u,c\},a\}+bu\{ \{ c,a\},u\}]+$$
$$  [cu\{ \{ a,b\},u\}+cu\{ \{ u,a\},b\}+cu\{ \{ b,u\},a\}]+$$
$$  [uu\{ \{ a,b\},c\}+uu\{ \{ b,c\},a\}+uu\{ \{ c,a\},b\}]=0.$$

The theorem is proved.

\

\subsection{Generalized Poisson algebras.}
The variety of unital {\it generalized Poisson algebras} (see, for example, \cite{kaysh14}) is
 defined by the identities
$$xy=yx, (xy)z=x(yz), \{xy,z\}=\{x,z\}y+x\{y,z\}+D(z)xy,D(x)=\{1,x\},$$
$$\{x,y\}=-\{y,x\}, \{\{x,y\},z\}+\{\{y,z\},x\}+\{\{z,x\},y\}=0.$$

    \begin{Th}Let $(A;\cdot,\{,\})$ be a generalized Poisson algebra.
    Then  $(A;[\{,\}, \cdot])$ is an associative-commutative algebra, and
    $(A;[\cdot, \{,\}])$ is a Lie  algebra.
    \end{Th}

{\bf Proof.}
In the first case, we have
$$a*b= \{u,ab\} - \{u,a\}b-a\{u,b\}=-D(u)ab=b*a$$
and
$$(a*b)*c=D(u)^2abc=a*(b*c).$$

In the second case,
$$a*b= u\{a,b\}-\{ua,b\}-\{a,ub\}=$$
$$- a\{u,b\}- b\{a,u\}-u\{a,b\} - D(b)ua+D(a)ub =-b*a.$$

Here we use the proof of Theorem \ref{ThPoisson}.
It is easy to see that
\begin{eqnarray*}
\circlearrowright_{a,b,c} [&& (a*b)*c]= \circlearrowright_{a,b,c} [(\{b,u\}a+\{b,a\}u+
\{u,a\}b-D(b)ua+D(a)ub)*c]=\\
\circlearrowright_{a,b,c} [&&\{c,u\}\{b,u\}a+\{c,b\}\{b,u\}a+\{c,\{b,u\}\}au-
\{b,u\}D(c)au+\{u,a\}\{b,u\}c+\\
&&\{u,\{b,u\}\}ac-\{b,u\}D(u)ac-\{b,u\}D(c)au+\{b,u\}D(a)cu+D(\{b,u\})acu+\\
&&\{c,u\}\{b,a\}u+\{c,u\}\{b,a\}u+\{c,\{b,a\}\}u^2-\{b,a\}D(c)u^2+\{u,\{b,a\}\}cu-\\
&& \{b,a\}D(u)cu-\{b,a\}D(c)u^2+\{b,a\}D(u)cu+D(\{b,a\})cu^2+\{c,u\}\{u,a\}b+\\
&& \{c,b\}\{u,a\}u+\{c,\{u,a\}\}bu-\{u,a\}D(c)bu+\{u,\{u,a\}\}bc+\{u,b\}\{u,a\}c-\\
&& \{u,a\}D(u)bc-\{u,a\}D(c)bu+\{u,a\}D(b)cu+D(\{u,a\})bcu-\{c,u\}D(b)au-\\
&& \{c,D(b)\}au^2-\{c,u\}D(b)ua-\{c,a\}D(b)u^2+2D(b)D(c)au^2-\{u,a\}D(b)cu-\\
&& \{u,D(b)\}acu+2D(b)D(u)acu+D(b)D(c)au^2 -D(D(b))acu^2-D(b)D(u)acu-\\
&& D(a)D(b)acu^2+\{c,u\}D(a)bu+\{c,D(a)\}bu^2+\{c,u\}D(a)bu+\{c,b\}D(a)u^2-\\
&& 2D(a)D(c)bu^2+\{u,D(a)\}bcu+\{u,b\}D(a)cu-2D(a)D(u)bcu-D(a)D(c)bu^2+\\
&& D(D(a))bcu^2+D(a)D(u)bcu+D(a)D(b)cu^2].\
\end{eqnarray*}

By the proof of Theorem \ref{ThPoisson}, we can conclude that the sum of all elements without $D$ is zero.
Now, we have
\begin{eqnarray*}
\circlearrowright_{a,b,c} [&& (a*b)*c]= \\
\circlearrowright_{a,b,c} [&&
-\{b,u\}D(c)au-\{b,u\}D(u)ac-\{b,u\}D(c)au+\{b,u\}D(a)cu+D(\{b,u\})acu-\\
&& \{b,a\}D(c)u^2-\{b,a\}D(u)cu-\{b,a\}D(c)u^2+\{b,a\}D(u)cu+D(\{b,a\})cu^2-\\
&& \{u,a\}D(c)bu-\{u,a\}D(u)bc-\{u,a\}D(c)bu+\{u,a\}D(b)cu+D(\{u,a\})bcu-\\
&& \{c,u\}D(b)au-\{c,D(b)\}au^2-\{c,u\}D(b)ua-\{c,a\}D(b)u^2+2D(b)D(c)au^2-\\
&& \{u,D(b)\}acu+2D(b)D(u)acu+D(b)D(c)au^2 -D(D(b))acu^2-D(b)D(u)acu-\\
&& \{u,a\}D(b)cu-D(a)D(b)acu^2+\{c,u\}D(a)bu+\{c,D(a)\}bu^2+\{c,u\}D(a)bu+\\
&& \{c,b\}D(a)u^2-2D(a)D(c)bu^2+\{u,D(a)\}bcu+\{u,b\}D(a)cu-2D(a)D(u)bcu-\\
&& D(a)D(c)bu^2+D(D(a))bcu^2+D(a)D(u)bcu+D(a)D(b)cu^2].\\
\end{eqnarray*}

It is easy to see that
\begin{eqnarray*}
\circlearrowright_{a,b,c} && [D(D(a))bcu^2-D(D(b))acu^2]=0,\\
\circlearrowright_{a,b,c} && [D(\{b,u\})acu+D(\{u,a\})bcu]=0,\\
\circlearrowright_{a,b,c} && [\{u,D(a)\}bcu-\{u,D(b)\}acu]=0,\\
\circlearrowright_{a,b,c} && [D(\{b,a\})cu^2-\{c,D(b)au^2+\{c,D(a)\}bu^2]=0,\\
\circlearrowright_{a,b,c} && [-2\{b,a\}D(c)u^2-\{c,a\}D(b)u^2+\{c,b\}D(a)u^2]=0,\\
\circlearrowright_{a,b,c} && [\{b,u\}D(c)au+\{u,a\}D(c)bu+\{c,u\}D(b)au+\{u,c\}D(a)bu]=0.
\end{eqnarray*}

Obviously, $\circlearrowright_{a,b,c} [(a*b)*c]=0$
and  $[\cdot, \{,\}]$ is a Lie  algebra.
The theorem is proved.

\

\subsection{Novikov-Poisson algebras.}
The variety of left {\it Novikov-Poisson} algebras is defined by the identities
$$xy=yx, (xy)z=x(yz),$$
$$x \circ (y \circ z)= y \circ (x \circ z), (x, y, z)_{\circ}  =  (x,z, y)_{\circ},$$
$$x\circ (y  z) = (x  \circ  y)  z,
(x   y) \circ  z - x   (y \circ  z)= (x  z) \circ y - x   (z \circ y).$$

    \begin{Th}Let $(A;\cdot,\circ)$  be a left Novikov-Poisson algebra.
    Then  $(A;[\cdot, \circ])$ is a left Novikov algebra and
          $(A;[\circ, \cdot])$ is an associative-commutative  algebra.
    \end{Th}

{\bf Proof.}
Firstly, we have
$$a * b = u (a \circ b)-(ua) \circ b-a \circ (ub)= -(ua)\circ b.$$
 Hence,
$$ a*(b*c)= (ua)\circ ((ub)\circ c)= (ub)\circ ((ua)\circ c)= b*(a*c),$$
and
$$(a,b,c)_*=(a*b)*c- a*(b*c)=
(u((ua)\circ b))\circ c - (ua)\circ ((ub)\circ c)=$$
$$((ua)\circ (ub))\circ c - (ua)\circ ((ub)\circ c) =
(ua, ub, c)_{\circ}=(ua, c, ub)_{\circ}=$$
$$=((ua)\circ c)\circ(ub)-(ua)\circ(c\circ(ub))=$$
$$u(ua,c,b)_{\circ}=u(ua,b,c)_{\circ}=(a*c)*b-a*(c*b)=(a,c,b)_*.$$
Secondly,
$$a * b = u \circ (ab)-(u\circ a) b-a(u\circ b)=-u\circ(ab)=b*a.$$
Therefore,
$$(a *  b) * c= u \circ ( (u \circ (ab)) c)=u \circ  (u \circ (abc))=u \circ (a (u \circ (bc)))= a*(b*c).$$
The theorem is proved.

\

Similarly, the variety of right Novikov-Poisson algebras may be defined  (see, for example, \cite{xu97}).
 It is easy to prove the following theorem:

    \begin{Th}Let $(A;\cdot,\circ)$  be a right Novikov-Poisson algebra.
    Then  $(A;[\cdot, \circ])$ is a right Novikov algebra and
          $(A;[\circ, \cdot])$ is a commutative  algebra.
    \end{Th}


\section{The Kantor square of algebras of special type.}

Here we study some special cases of the Kantor square.
For an algebra $A:=(A;\cdot)$  its the Kantor square $(A;[\cdot,\cdot])$ we denote by $(A,*).$
We denote the Kantor square for a fixed element $u$ by $(A,*_u)$.
We consider the relations between the ideals in $A$ and $(A,*)$,
the relations between an associative algebra $A$ with polynomial identity  and its the Kantor square.
Moreover, the relations between the nilpotency and right nilpotency in $A$ and $(A,*)$ are  investigated.

\subsection{Ideals in the Kantor product.}

\begin{Th}\label{}
Let $I$ be an ideal of  $A$. Then $(I,*)$ is an ideal of $(A,*),$
but the converse statement  is not true in general.
\end{Th}

{\bf Proof.}
It is easy to see that
if $i \in I$ and $a \in A$ then
$$i*a=u(ia)-(ui)a-i(ua) \in I \mbox{ and }a*i=u(ai)-(ua)i-a(ui) \in I.$$
It  follows that $I$ is an ideal of $(A,*).$

Conversely,
we can consider the trivial case,
where for an algebra $A$ has zero Kantor square (for example, Lie or Leibniz algebra)
and every subspace of $A$ is an ideal of $(A,*).$
For the non-trivial case (nonzero Kantor product),
we can consider the following associative algebra: $A_1 \oplus A_2$
 is the direct sum of the matrix algebras of order $2$.
 Here, if $e_i$ is the unit of $A_i$ then
 the subspace generated by $A_1$ and $e_2$ is an ideal of $(A,*_{e_1}),$
but is not an ideal of $A_1 \oplus A_2.$
The theorem is proved.

\subsection{Associative algebras with polynomial identity.}
Given a polynomial $f$ in $n$ variables,
we define $f_*(x_1, \ldots, x_n)$
as the value of $f$ in $(A,*)$, where $x_1, \ldots, x_n$ are some elements in $ A.$

\begin{Th}\label{ass_id1}
Let $(A;\cdot)$ be an associative algebra that satisfies the polynomial identity $f(x_1,\ldots, x_n)$.
 Then there exists an identity $g$ such that
 $A$ and $(A,*)$ satisfy $g$.
\end{Th}

{\bf Proof.}
It is easy to see
that if $A$ satisfies the identity $f(x_1, \ldots, x_n)$
then $A$ satisfies the identity $g(x_1, \ldots, x_n,z)=f(x_1, \ldots, x_n)z.$
By Theorem \ref{thass},
the multiplication in algebra $(A,*)$ is defined by $x*y=-xuy.$
Now, we can calculate the element $g_*(x_1, \ldots, x_n,z)$ in $(A,*)$.
Obviously, it is $(-1)^{n}f(x_1u, \ldots, x_nu)z$ which amounts to zero in $A$.
It follows that $(A,*)$ satisfies the identity $g.$  
The theorem is proved.

One of the most popular identity in the associative algebras is the standard polynomial identity of degree $n$:
$$s_{n}(x_1, \ldots, x_n)= \sum_{\sigma \in S_{n}} (-1)^{\sigma} x_{\sigma(1)} \ldots  x_{\sigma(n)}.$$

\begin{Th}\label{ass_id2}
Let $(A;\cdot)$ be an associative algebra that
 satisfies $s_n$. Then $(A,*)$ satisfies $s_{n+1}.$

\end{Th}

{\bf Proof.}
It is easy to see that
 the standard polynomial of degree $n+1$ may be written as
$$ s_{n+1}(x_1, \ldots, x_{n+1})=
\sum_{i=1}^n (\sum_{ \begin{array}{c} \small\sigma \in S_{n+1}, \\ \sigma(n+1)=i
\end{array}} (-1)^{\sigma} x_{\sigma(1)} \ldots  x_{\sigma(n)} x_{\sigma(n+1)=i})=$$
$$=\sum_{i=1}^n (  \epsilon_{i} s_{n}(x_1, \ldots, \hat{x_i}, \ldots, x_{n+1}) x_{i}),
\epsilon_{i} = \pm 1.$$
Now, by the proof of Theorem \ref{ass_id1},  $(A,*)$ satisfies
 the standard polynomial identity of degree $n+1.$
The theorem is proved.

\subsection{Nilpotent algebras.}
For the nilpotent algebras, we can prove the following theorem.

\begin{Lem}\label{nilpotent}
Let $(A;\cdot)$ be a nilpotent algebra of nilpotency index $n$.
Then $(A,*)$ is a nilpotent algebra of nilpotency index $\leq [n/2]+1.$
\end{Lem}

{\bf Proof.}
Obviously, every product of the form  $x_1*x_2*\ldots *x_t$ (with some order of brackets) 
is a sum of
 multiplications of the form $y_1y_2\cdots y_{2t-1}$ (with some order of brackets).
 Now, it is easy to see that $(A,*)$ is nilpotent and its index of nilpotency $\leq [n/2]+1.$
The Lemma is proved.

\subsection{Right-nilpotent algebras.}
An algebra $A$ is called right-nilpotent (or left-nilpotent) of nilpotency index $n$
 if it satisfies the identity
$$(\ldots (x_1x_2)\ldots)x_n=0 \ \ \ (\mbox{ or }   x_1( \ldots (x_{n-1}x_n)\ldots )=0).$$
Curiously, an analogue of the Theorem \ref{nilpotent} is not true for the right-nilpotent algebras.

\begin{Th}\label{rightnilpotent}
There exists a right nilpotent algebra
 $(A;\cdot)$  such that $(A,*_u)$ is not right nilpotent, but  $(A,*_u)$  is solvable.
\end{Th}

{\bf Proof.} An algebra $A$ is \textit{right alternative} if the following identity holds in $A$:
\begin{equation*} (x,y,y) = 0. 
\end{equation*}
It is interesting fact that in contrast to the algebras of many well-studied classes
 (Jordan, alternative, Lie and so on) a right nilpotent right alternative algebra need not be non-nilpotent.
The corresponding example of a five-dimensional right nilpotent but not nilpotent algebra belongs to
  Dorofeev \cite{Dorofeev}.
Its basis is $\{a, b, c , d , e \},$  and the multiplication is given
 by (zero products of basis vectors are omitted)
$$ab = -ba = ae = -ea = db = -bd = -c, ac = d, bc = e.$$
It is easy to see that
$$c *_ab = a(cb)-(ac)b-c(ab)=c.$$
Obviously, $c=( \ldots (c*_a b)*_a \ldots )*_ab \neq 0,$
and $(A,*_a)$ is not right-nilpotent.
It is easy to see that $A^2 \subseteq \langle c,d,e \rangle$ and $A*_aA  \subseteq \langle c,d,e \rangle$,
 but $(A*_aA)*_a(A*_aA)=0,$
 and $(A,*_a)$ is solvable.

The theorem is proved.

\subsection{Derivations.}
Remember that a linear mapping $D$ of an algebra $A$ is called a derivation
if it satisfies the relation $D(xy)=D(x)y+xD(y).$
By \cite{Kantor90}, an element $a$ of an algebra $A$ is called a {\it Jacobi} element if
if satisfies the relation $a(xy)=(ax)y+x(ay).$ 
All Jacobi elements of $A$ form a vector space, which is  called the {\it Jacobi} space of $A.$

\begin{Lem}\label{derivation}
Let $D$ be a derivation of both $A$ and $(A,*).$
Then

$1)$ If $A$ has zero Jacobi space, then $D=0;$

$2)$ If $D$ is invertible, then $A$ is a left Leibniz algebra and $(A,*)$ is a zero algebra. 
In particularly, if $A$ is a finite-dimensional algebra over a field of zero characteristic, then $A$ is nilpotent.

\end{Lem}

{\bf Proof.} 
$1).$ By simple calculations, from
$D(x*y)=D(x)*y+x*D(y),$
we have  
$$D(u)(xy)=(D(u)x)y+x(D(u)y).$$
By the definition of the Jacobi space, we have $D=0.$

$2).$ By invertibility of mapping $D$ and arbitrarity of element $u,$
we infer that $A$ is a left Leibniz algebra.
By the Lemma \ref{leibniz}, we imply that $(A,*)$ is zero algebra.

In \cite{Fialc12} it was proved that a finite-dimensional Leibniz algebra
over a field of characteristic zero which admitting an invertible derivation is nilpotent.
The Lemma is proved.

\subsection{Automorphisms.}
Remember that an invertible linear mapping $\phi$ of an algebra is called an automorphism
if it satisfies the relation $\phi(xy)=\phi(x)\phi(y).$

\begin{Lem}\label{derivation}
Let $\phi$ be an automorphism of both $A$ and $(A,*).$
If $A$ is an algebra with zero Jacobi space, then $\phi$ is the identity mapping.
\end{Lem}

{\bf Proof.} 
By simple  calculations from
$\phi(x*y)=\phi(x)*\phi(y),$
we have  
$$(u-\phi(u))(xy)=((u-\phi(u))x)y+x((u-\phi(u))y).$$
By the definition of the Jacobi space, we have that $\phi=id.$
The Lemma is proved.

\subsection{Isomorphic Kantor squares.}
Here we talk about the situation where  algebra $A$ and its Kantor square are isomorphic.

\begin{Th}\label{Isomor}
Let $A$ be a finite-dimensional associative algebra.
Then $A$ is isomorphic to $(A,*),$
if and only if $A$ is a skew field.

\end{Th}

{\bf Proof.} 
Let $f_u$ is an isomorphism between algebras $A$ and $(A,*_u)$
and $f_u(xy)=f_u(x)*f_u(y)=-f_u(x)uf_u(y).$
If in $A$ there are two elements $u$ and $v$ with zero product,
for $x=f_u^{-1}(v)$ 
we have 
$$f_u(f_x(ab))=-f_u(f_x(a)xf_x(b))=f_u(f_x(a))uvuf_u(f_x(b))=0.$$
Now, if there is a zero divisor, then the algebra $A$ has zero multiplication.
It is Well known that every finite-dimensional algebra without zero divisors is a skew field.

On the other side, for some fixed nonzero element $u$ from a skew field $A$ 
we define $f_u(a)=-au^{-1}.$
It is an isomorphism between algebras $A$ and $(A,*_u)$ for every nonzero element $u.$
The theorem is proved.

\

{\bf Acknowledgments.}
I am grateful to Prof. Dr. Ivan Shestakov for the idea of this work,
Prof. Dr. Alexandre Pojidaev and Yury Popov for some correction of the language,
Dr. Yury Volkov for the idea of the proof of the Theorem \ref{Isomor}.


    \end{document}